OpenStudy (anonymous):

If RS = 2x + 5, RU = x + 8, and TU = x + 16, what is the value of x?

OpenStudy (anonymous):

is there a diagram with this q?

OpenStudy (anonymous):



OpenStudy (anonymous):

is RSTU a parallelogram

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok - so that means opposite sides are equal so RS = TU and 2x + 5 = x + 16 solving for x: 2x - x = 16 - 5 x = 11

OpenStudy (anonymous):

thankyou

OpenStudy (anonymous):

yw